OpenStudy (anonymous):

find the slope of the line by solving for y x - 3y = -2 please show all steps

OpenStudy (anonymous):

y=2/3+1/3x so ur slope is 1/3

OpenStudy (saifoo.khan):

SHOW STEPS.

OpenStudy (anonymous):

how did you get y=2/3+1/3x

OpenStudy (anonymous):

x + 2 = 3y (x+2)/3 = y y = (1/3)x + 2/3

OpenStudy (anonymous):

x - 3y = -2 -x -x -3y = -x -2 /-3 /-3 y = 1/3x + 2/3

OpenStudy (anonymous):

u have to solve for Y and the factor of X will be ur slope of the line ! i moved x to the other side of the equation first then divided everything by a negative 3 and that's the result u end up with